OpenStudy (anonymous):

Determine if the series abs converges, conditionally converges, or diverges. n^n /n!

OpenStudy (anonymous):

As \( n\to \infty\) \[n!>>n^2\] so the series converges

OpenStudy (anonymous):

Using ratio test, I found that it diverges.. But I don't know what test to use to find if it conditionally converges. :/

OpenStudy (anonymous):

If it diverges, you don't have to worry about conditional convergence. In general, though, you'd have to show that a given series isn't absolutely convergent. For this, you'd have to show that the following isn't true: \(\displaystyle\sum |a_n|\text{ converges.}\)

OpenStudy (anonymous):

For this series, its always positive. I'm not sure why I used ratio test but I did. I'm confused at when I can use ratio and root tests (only when it's alternating signs?). And I thought a series can converge even if it absolutely diverges

OpenStudy (anonymous):

"For this series, it's always positive." Yes. "I'm not sure why I used the ratio test..." If your series contains a factorial, the ratio test usually works pretty well. "...confused when to use ratio/root tests (only when alternating signs?)..." Here's a handy chart for knowing when to use which test: http://www.math.tamu.edu/~austin/serieschart.pdf As for the alternating sign - use the alternating series test. The reason why should be fairly clear. "I thought a series can converge even if it absolutely diverges." That's not at all the case. If a series diverges, it diverges. It can't converge if it diverges. If a series converges, then it can be either conditionally or absolutely convergent.